ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

14. Mathematical Aspects of Computer Science

1991 MS Classification: 18, 68 Ehrig, Hartmut, Technical University of Berlin, Germany On the Role of Category Theory in Computer Science The paradigm and the main concepts of category theory are summarized in view of application in computer science. Some areas of computer science influenced by category theory are addressed. Especially the role of category theory in the area of algebraic specifications, graph transformation and Petri nets, is demonstrated.